# Human Action Recognition from Point Clouds over Time

### v1

@@ I Introduction
With the proliferation of consumer grade depth sensors such as Microsoft’s Kinect series (v1, v2, Azure), as well as Lidar-endowed robots and instruments, researchers have begun to develop methods for human action recognition based on three dimensional video data, sampling a fixed number of points per frame as input, discussed in Section II-B. These approaches largely draw from the existing modern 3D computer vision literature concerning point cloud classification, segmentation, and object detection. Existing methods do not segment the individuals from the background beyond simple depth thresholding or histogram techniques, and rely purely on point-based approaches that require extensive local spatio-temporal neighborhood grouping computations to group points in both space and time. Further, they require a depth sensor or multi-view camera setups to obtain depth information, which is not practical or available in many settings. Recent research into monocular depth estimation [16, 17], providing the ability to estimate metric or ordinal disparity maps from images and videos, now offers the promise of estimated point clouds, which we explore in this work.
@@
  • Depth sensing technology is present, in which case we have real world point clouds sequences.
  • An RGB video is used as input, in which case we propose to use modern monocular depth estimation to obtain estimated point clouds.
@@
Inspired by recent research into 3D object detection based on voxelized point clouds, alongside point-based approaches, we develop a novel backbone for human action recognition. Our approach integrates sparse convolution and point-wise embeddings. The input to this model consists of segmented point clouds, which are obtained through our proposed pipeline.
@@ II-A Point Cloud Deep Learning
As noted by Lu et al. [18], the focus of modern point cloud deep learning has largely been model classification, part segmentation, semantic segmentation, and 3D object detection and tracking. Other topics of interest include point cloud registration, learnable downsampling and upsampling, denoising, and scene completion from occluded inputs. Point cloud methods can be distinguished by two main tracks. The first track are those approaches that employ purely point-based operations to learn features from localized aggregation operators. Neighborhoods are estimated using K-nearest neighbors (K-nn), ball radius queries, or windows of points serialized with space-filling curves. A second track maps points to 3D grid locations, known as voxels, and uses sparse convolution or attention operations based on non-empty regions to compute features. In this work we do not focus on point cloud models that employ projection to images, but rather leave the original geometry largely intact. In this section we briefly summarize representative works in both tracks.
@@ II-A1 Point-Based Models
In keeping with the trend in modern computer vision to explore the use of concepts from the Transformer architecture [22], PointTransformer v1 [23] makes use of self-attention mechanisms to aggregate features in local radii around points, using relative positional embeddings, employed for both point cloud classification and segmentation. The follow-up model PointTransformer v2 [24] utilizes channel groupings in the attention mechanism, and uniform grid-based pooling in the down-sampling and upsampling stages of the network. The most recent PointTransformer v3 [25] focuses on large-scale semantic segmentation by using point serialization with space-filling curves (z-order curves and Hilbert curves), to learn features within windows of 1 dimensional arrays of points, introducing serialized pooling/unpooling. They leverage serialization to avoid memory-expensive neighborhood searches. In a similar vein, Wang develops the OctFormer [26] model, which also incorporates z-order curves to reshape points into windows, computing attention within windows, but computes conditional positional encoding of points using octree-based convolution.
@@ II-B Action Recognition from Point Cloud Videos
Distinct from the approach proposed in this paper, the preceding models do not provide human segmentation in image space or 3D space, leaving a large number of irrelevant background points in the data, and moreover do not segment body parts, or use surface normals. Further, they all consider that depth is measured from a sensor, but never estimated with monocular depth estimation. The models detailed in the previous literature review do not explore hybrid sparse convolution and point-based approaches. We aim to explore these alternatives.
@@ III-A Human Point Clouds From Aligned Depth/IR Sequences
In order to obtain human point clouds over time, we first consider the scenario where we have aligned pairs of equal-resolution images $(I_{t},D_{t})_{t=1}^{T}$ for a sequence of $T$ frames, where $D_{t}$ are depth images. In the case of RGB-D images, $I_{t}$ will be an RGB image. With respect to experiments conducted in Section IV, since the NTU RGB-D dataset does not provide aligned RGB/depth images, rather we opt to consider $I_{t}$ as the pixel aligned infrared (IR) images. It is assumed that camera intrinsics are available. An overview of the proposed pipeline is shown in Figure 1.
@@
Initially, simultaneous instance and body part segmentation, often referred to as human parsing, is performed on each individual image $I_{t}$ in a sequence. For this, we opt to use the high performing model M2FP (Mask2FormerParsing) [38], based off the Mask2Former architecture [39], which computes instance, body part, and background masks from learnable mask queries input to a Transformer decoder. Instance masks are then denoised with simple heuristics, including removing very small connected components, and filling in the convex hull of instance masks to remove small interior holes caused by parsing inaccuracy. The result for frame $t$ is a set of $N$ instance masks $M_{t,n}\in\{0,1\}^{H\times W}$ and per-image body part masks $P_{t}\in\{0,C+1\}^{H\times W}$ for $C$ non-background classes, with 0 given to the background label.
@@ III-B Human Point Clouds from RGB using Monocular Depth Estimation
It is often the case that the input to an action recognition algorithm is an RGB video from a single view, with no depth sensing technology present. Hence, capturing 3D geometry in this scenario is challenging. However, recent development of algorithms for monocular depth estimation has seen remarkable progress, where for a given color image, a disparity map is estimated. Modern approaches often train on mixtures of real-world datasets obtained with depth sensors, and photorealistic synthetic images rendered in computer graphics engines, focusing on fine-grained structural details as well as ordinal relations between points. In this work, we opt for the Depth-Anything v2 (DAv2) approach for its fine-grained accuracy, crucial for action recognition [17]. An overview of the proposed pipeline in this setting is shown in Figure 2.
@@
where $\epsilon$ is a small constant used to prevent division by zero. Denoising of instance masks, in addition to denoising in 3D is similar to section III-A, using DBSCAN and a percentile-based denoising strategy rather than a metric distance. In particular, points that are outside percentile range of 5905-90 of distance from the projected instance mask centroid are pruned. Since RGB images often have much higher resolution that depth images, we apply sampling to windows of points sorted by serialization with space-filling curves [25], reducing the complexity for $K$ sample points, $W$ windows, and $N$ overall points from $O(NK)$ to $O(NK/W)$. Surface normals are also estimated in this setting.
@@ III-C Action Recognition From Human Point Clouds over Time
In this section, we detail the proposed action recognition model. We now assume that, given the denoised output from sections III, the input to the model are batches of tensors $S$ of shape $(B,T,N,C)$, for $T$ input frames, $N$ points, and $C$ channels, and batch size $B$. In the setting of point cloud sequences obtained with depth sensors, we consider channels corresponding to $3D$ coordinates, surface normals, infrared intensity, and part label features. Part labels are converted into learnable embeddings of dimension 3 using an embedding table. In the setting of point clouds estimated with monocular depth estimation, identical features are used, replacing infrared with RGB values. An overview is shown in Figure 3.
@@ III-C1 T-Net Embeddings
The output $T_{\text{final}}\in\mathbb{R}^{B^{\prime}\times C\times C}$ is used to transform the input point cloud using batch matrix multiplication, after which we concatenate the original features to retain both local and global information, i.e.
@@ III-C2 Sparse CNN Backbone
Before presenting the proposed sparse CNN backbone, we first review sub-manifold sparse convolution [30], in 3D for completeness. We opt for the sub-manifold variant due to its computational efficiency, reducing the dilution effect of regular sparse convolution, which we restrict to sparse max pooling operations. For a grid size $(g_{x},g_{y},g_{z})$, each point $p=(p_{x},p_{y},p_{z})$ in each frame is mapped to a voxel location $(x^{\prime},y^{\prime},z^{\prime})\in\mathbb{Z}_{+}^{3}$ by coordinate division by the grid size, rounding down.
@@
where $O(K)$ are the offset location of kernel $K$ with non-empty voxels relative to center location $(x^{\prime},y^{\prime},z^{\prime})$, and $A$ is an aggregation operation for all points falling into a given voxel for channel $c$, typically the mean or max value. For all sparse convolutions in the following, we assume the sub-manifold formulation, with no dilation, and zero-padding voxels to preserve the input resolution.
The proposed backbone is detailed in Table 1, with layer descriptions to follow, where we refer to the corresponding architecture as SP-HP-ConvoT (Sparse Convolution for Human Point Clouds over Time). Note that in Table 1, and Figure 4, output shapes are shown in dense voxelized form, but the sparse convolution data structure stores features per point, and their respective spatio-temporal grid and batch indices.
@@ IV-C Results
In the following, we refer to SP-HP-ConvoT for models using point cloud sequences from a depth sensor, and SP-HP-ConvoT-MDE for point cloud sequences obtained from monocular depth estimation (MDE). In Table 2 we observe the effects of the different auxiliary channels considered, namely infrared, surface normals, part label embeddings, and color.
@@
Next, in Table 4, we compare our model against three architectures from the skeletal action recognition literature (SAR). We focus on models that take either the raw keypoint representation, commonly referred to as joints, or the bones representation as input. The bones input format consists of link vectors between keypoints, defined by an adjacency matrix, and is frequently used in joint-bone ensembles, where predictions from separate models are combined with weighted contributions. The proposed approach is competitive, lightly outperforming existing reported methods on the CSet split.
@@
Inference speed is evaluated in Table 5 based on the number of points sampled from the point cloud and people, measured in sequences per second on a single Nvidia RTX 4090 GPU. The evaluation uses sequence lengths of 32 frames with a batch size of 1, considering only raw geometry (i.e., 3 input channels). For comparison, we include the SAR baseline ST-GCN++ [43] using 133 keypoints from the COCO WholeBody dataset [49], selected for its relatively dense keypoint representation, with either one or two people. While the proposed approach runs slower, it achieves 40 sequences per second with 2048 points when processing two people.
@@
In Table 6, we compare the results of existing point cloud action recognition models to the proposed approach, where we also consider an ensemble of the SP-HP-ConvoT and SP-HP-Convot-MDE models, using weights $\lambda_{1},\lambda_{2}$ obtained through optimizing performance on a separate validation set during training, i.e. predictions of the form
@@
On the CSub split, the SP-HP-ConvoT model with auxiliary features outperforms existing approaches by 1.0%. The proposed ensemble achieves the best result (+2.3%). However, both in the single model and ensemble setting on the CSet benchmark, performance lags behind. We hypothesize that the accuracy loss may stem from the cumulative effect of person and body part segmentation inaccuracies. If the model has not encountered similar noise patterns or collection setups during training, which vary in the CSet setting, it may struggle to distinguish body points from the background as effectively.
@@ V Conclusion
In conclusion, this work presents a novel approach to human action recognition from point cloud sequences. We employed a pipeline for segmenting body parts and tracking individuals, leveraging 3D projection with depth images and, alternatively, monocular depth estimation. A novel backbone using sparse convolutional neural networks was developed, which to our knowledge is the first of its kind in the literature for human action recognition from point clouds. Model performance was validated on a popular action recognition dataset, showing competitive results and a new benchmark high for temporal point cloud deep learning on the NTU RGB-D 120 dataset within the CSub split.

### v2

@@ I Introduction
With the proliferation of consumer grade depth sensors such as Microsoft’s Kinect series (v1, v2, Azure), as well as Lidar-endowed robots and instruments, researchers have begun to develop methods for human action recognition based on three dimensional video data, sampling a fixed number of points per frame as input, discussed in Section II-B. These approaches largely draw from the existing modern 3D computer vision literature concerning point cloud classification, segmentation, and object detection. Existing methods do not segment the individuals from the background beyond simple depth thresholding or histogram techniques, and rely purely on point-based approaches that require extensive local spatio-temporal neighborhood grouping computations to group points in both space and time. Further, they require a depth sensor or multi-view camera setups to obtain depth information, which is not practical or available in many settings. Recent research into monocular depth estimation [16, 17], providing the ability to estimate metric or ordinal disparity maps from images and videos, now offers the promise of estimated point clouds, which are explored in this work.
@@
  • Depth sensing technology is present, in which case real world point clouds sequences are considered.
  • An RGB video is used as input, in which case modern monocular depth estimation algorithms are used to obtain estimated point clouds.
@@
Inspired by recent research into 3D object detection based on voxelized point clouds, alongside point-based approaches, a novel backbone for human action recognition is developed. Our approach integrates sparse convolution and point-wise embeddings. The input to this model consists of segmented point clouds, which are obtained through our proposed pipeline.
@@ II-A Point Cloud Deep Learning
As noted by Lu et al. [18], the focus of modern point cloud deep learning has largely been model classification, part segmentation, semantic segmentation, and 3D object detection and tracking. Other topics of interest include point cloud registration, learnable downsampling and upsampling, denoising, and scene completion from occluded inputs. Point cloud methods can be distinguished by two main tracks. The first track are those approaches that employ purely point-based operations to learn features from localized aggregation operators. Neighborhoods are estimated using K-nearest neighbors (K-nn), ball radius queries, or windows of points serialized with space-filling curves. A second track maps points to 3D grid locations, known as voxels, and uses sparse convolution or attention operations based on non-empty regions to compute features. In this section, point cloud models that employ projection to images are not considered, but rather representative works in both tracks are summarized.
@@ II-A1 Point-Based Models
In keeping with the trend in modern computer vision to explore the use of concepts from the Transformer architecture [22], PointTransformer v1 [23] makes use of self-attention mechanisms to aggregate features in local radii around points, using relative positional embeddings, employed for both point cloud classification and segmentation. The follow-up model PointTransformer v2 [24] utilizes channel groupings in the attention mechanism, and uniform grid-based pooling in the down-sampling and upsampling stages of the network. The most recent PointTransformer v3 [25] focuses on large-scale semantic segmentation by using point serialization with space-filling curves (z-order curves and Hilbert curves), to learn features within windows of 1 dimensional arrays of points, introducing serialized pooling/unpooling. They leverage serialization to avoid memory-expensive neighborhood searches. In a similar vein, Wang develops the OctFormer model [26], which also incorporates z-order curves to reshape points into windows, computing attention within windows, but computes conditional positional encoding of points using octree-based convolution.
@@ II-B Action Recognition from Point Cloud Videos
Distinct from the approach proposed in this paper, the preceding models do not provide human segmentation in image space or 3D space, leaving a large number of irrelevant background points in the data, and moreover do not segment body parts, or use surface normals. Further, they all consider that depth is measured from a sensor, but never estimated with monocular depth estimation. The models detailed in the previous literature review do not explore hybrid sparse convolution and point-based approaches, hence the motivation to explore these alternatives.
@@ III-A Human Point Clouds From Aligned Depth/IR Sequences
In order to obtain human point clouds over time, first consider the scenario of aligned pairs of equal-resolution images $(I_{t},D_{t})_{t=1}^{T}$ for a sequence of $T$ frames, where $D_{t}$ are depth images. In the case of RGB-D images, $I_{t}$ will be an RGB image. With respect to experiments conducted in Section IV, since the NTU RGB-D dataset does not provide aligned RGB/depth images, treat $I_{t}$ as the pixel aligned infrared (IR) images. It is assumed that camera intrinsics are available. An overview of the proposed pipeline is shown in Figure 1.
@@
Initially, simultaneous instance and body part segmentation, often referred to as human parsing, is performed on each individual image $I_{t}$ in a sequence. For this, the high performing model M2FP (Mask2FormerParsing) [38] was employed, based off the Mask2Former architecture [39], which computes instance, body part, and background masks from learnable mask queries input to a Transformer decoder. Instance masks are then denoised with simple heuristics, including removing very small connected components, and filling in the convex hull of instance masks to remove small interior holes caused by parsing inaccuracy. The result for frame $t$ is a set of $N$ instance masks $M_{t,n}\in\{0,1\}^{H\times W}$ and per-image body part masks $P_{t}\in\{0,C+1\}^{H\times W}$ for $C$ non-background classes, with 0 given to the background label.
@@ III-B Human Point Clouds from RGB using Monocular Depth Estimation
It is often the case that the input to an action recognition algorithm is an RGB video from a single view, with no depth sensing technology present. Hence, capturing 3D geometry in this scenario is challenging. However, recent development of algorithms for monocular depth estimation has seen remarkable progress, where for a given color image, a disparity map is estimated. Modern approaches often train on mixtures of real-world datasets obtained with depth sensors, and photorealistic synthetic images rendered in computer graphics engines, focusing on fine-grained structural details as well as ordinal relations between points. In this work, the Depth-Anything v2 (DAv2) approach was chosen for its fine-grained accuracy, crucial for action recognition [17]. An overview of the proposed pipeline in this setting is shown in Figure 2.
@@
where $\epsilon$ is a small constant used to prevent division by zero. Denoising of instance masks, in addition to denoising in 3D is similar to section III-A, using DBSCAN and a percentile-based denoising strategy rather than a metric distance. In particular, points that are outside percentile range of 5905-90 of distance from the projected instance mask centroid are pruned. Since RGB images often have much higher resolution that depth images, sampling was applied to windows of points sorted by serialization with space-filling curves [25], reducing the complexity for $K$ sample points, $W$ windows, and $N$ overall points from $O(NK)$ to $O(NK/W)$. Surface normals are also estimated in this setting.
@@ III-C Action Recognition From Human Point Clouds over Time
In this section, the proposed action recognition model is detailed. It is now assumed that, given the denoised output from sections III, the input to the model are batches of tensors $S$ of shape $(B,T,N,C)$, for $T$ input frames, $N$ points, and $C$ channels, and batch size $B$. In the setting of point cloud sequences obtained with depth sensors, channels consist of $3D$ coordinates, and possible auxiliary features, namely surface normals, infrared intensity, and part label features. Part labels are converted into learnable embeddings of dimension 3 using an embedding table. In the setting of point clouds estimated with monocular depth estimation, identical features are used, replacing infrared with RGB values. An overview is shown in Figure 3.
@@ III-C1 T-Net Embeddings
The output $T_{\text{final}}\in\mathbb{R}^{B^{\prime}\times C\times C}$ is used to transform the input point cloud using batch matrix multiplication, after which the original features are concatenated to retain both local and global information, i.e.
@@ III-C2 Sparse CNN Backbone
Before presenting the proposed sparse CNN backbone, the sub-manifold sparse convolution [30], in 3D, is reviewed for completeness. The sub-manifold variant was chosen due to its computational efficiency, reducing the dilution effect of regular sparse convolution, restricting to sparse max pooling operations. For a grid size $(g_{x},g_{y},g_{z})$, each point $p=(p_{x},p_{y},p_{z})$ in each frame is mapped to a voxel location $(x^{\prime},y^{\prime},z^{\prime})\in\mathbb{Z}_{+}^{3}$ by coordinate division by the grid size, rounding down.
@@
where $O(K)$ are the offset location of kernel $K$ with non-empty voxels relative to center location $(x^{\prime},y^{\prime},z^{\prime})$, and $A$ is an aggregation operation for all points falling into a given voxel for channel $c$, typically the mean or max value. For all sparse convolutions in the following, the assume the sub-manifold formulation is assumed, with no dilation, and zero-padding voxels to preserve the input resolution.
The proposed backbone is detailed in Table 1, with layer descriptions to follow, where the corresponding architecture will be referred to as SP-HP-ConvoT (Sparse Convolution for Human Point Clouds over Time). Note that in Table 1, and Figure 4, output shapes are shown in dense voxelized form, but the sparse convolution data structure stores features per point, and their respective spatio-temporal grid and batch indices.
@@ IV-C Results
In the following, SP-HP-ConvoT refer to models using point cloud sequences from a depth sensor, and SP-HP-ConvoT-MDE for point cloud sequences obtained from monocular depth estimation (MDE). In Table 2, the effects of the different auxiliary channels are considered, namely infrared, surface normals, part label embeddings, and color.
@@
Next, in Table 4, the proposed model is compared against three architectures from the skeletal action recognition literature (SAR). In this comparison comparisons were made from models that take either the raw keypoint representation, commonly referred to as joints, or the bones representation as input. The bones input format consists of link vectors between keypoints, defined by an adjacency matrix, and is frequently used in joint-bone ensembles, where predictions from separate models are combined with weighted contributions. The proposed approach is competitive, lightly outperforming existing reported methods on the CSet split.
@@
Inference speed is evaluated in Table 5 based on the number of points sampled from the point cloud and people, measured in sequences per second on a single Nvidia RTX 4090 GPU. The evaluation uses sequence lengths of 32 frames with a batch size of 1, considering only raw geometry (i.e., 3 input channels). For comparison, the SAR baseline ST-GCN++ [43] is included, using 133 keypoints from the COCO WholeBody dataset [49], selected for its relatively dense keypoint representation, with either one or two people. While the proposed approach runs slower, it achieves 40 sequences per second with 2048 points when processing two people.
@@
In Table 6, the results of existing point cloud action recognition models to the proposed approach is compared, also considering an ensemble of the SP-HP-ConvoT and SP-HP-Convot-MDE models, using weights $\lambda_{1},\lambda_{2}$ obtained through optimizing performance on a separate validation set during training, i.e. predictions of the form
@@
On the CSub split, the SP-HP-ConvoT model with auxiliary features outperforms existing approaches by 1.0%. The proposed ensemble achieves the best result (+2.3%). However, both in the single model and ensemble setting on the CSet benchmark, performance lags behind. It may be that the accuracy loss may stem from the cumulative effect of person and body part segmentation inaccuracies. If the model has not encountered similar noise patterns or collection setups during training, which vary in the CSet setting, it may struggle to distinguish body points from the background as effectively.
@@ V Conclusion
In conclusion, this work presents a novel approach to human action recognition from point cloud sequences. A pipeline for segmenting body parts and tracking individuals was employed, leveraging 3D projection with depth images and, alternatively, monocular depth estimation. A novel backbone using sparse convolutional neural networks was developed, which to our knowledge is the first of its kind in the literature for human action recognition from point clouds. Model performance was validated on a popular action recognition dataset, showing competitive results and a new benchmark high for temporal point cloud deep learning on the NTU RGB-D 120 dataset within the CSub split.